OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? A.) {2, 3,√13 } B.) {2, 2, 4} C.) (1, 2, √3}

OpenStudy (welshfella):

check if the square of longest side = sum of squares on the other 2 sides (Pythagoras:- c^2 = a^2 + b^2)

OpenStudy (welshfella):

is (√13)^2 = 2^2 + 3^2 ?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

The answer is A. Use Pythagorean Theorem: 2^2 + 3^2 = \sqrt{13}^2

OpenStudy (anonymous):

\[2^2 + 3^2 + \sqrt{13}^2\]